Temperature of a room at midnight is -4°C. The room temperature increases at the rate of 2°C per hour. Find the temperature at 4am and 12 noon.
1
Expert's answer
2017-06-07T10:53:12-0400
T(t)=-4+2t.
At 4 am T(4)=-4+2*4=4^0 C.
At 12 noon T(12)=-4+2*12=〖20〗^0 C.